Selenium in a 10.0g soil sample is distilled as the tetrabromide, which is collected in aqueous solution where it is hydrolyzed to SeO32-. The SeO32-​ is determined iodometrically requiring 4.5ml of standard thiosulfate solution for the titration. If the theosulfate titer is 0.049mg K2Cr2O7/ml, what is the concentration of selenium in the soil in ppm?

Thanks!
Answer #1

The reactions taking place are:

Se + 2 Br2 ------> SeBr4 …..(1)

2 SeBr4 + 3 H2O -------> Se2O32- + 6 H+ + 8 Br- …..(2)

Se2O32- + 4 I- + 6 H+ ------> 2 Se + I2 + 3 H2O ……(3)

I2 + 2 S2O32- -------> 2 I- + S4O62- …..(4)

As per the balanced stoichiometric reaction (4),

1 mole I2 = 2 moles thiosulfate

Again, the standardization of thiosulfate with dichromate can be shown as:

6 I- + Cr2O72- + 14 H+ -------> 3 I2 + 2 Cr3+ + 7 H2O ……(5)

I2 + 2 S2O32- -------> 2 I- + S4O62- ……(6)

As per equations (5) and (6),

1 mole Cr2O72- = 3 moles I2

1 mole I2 = 2 moles thiosulfate.

Therefore, 1 mole Cr2O72- = 6 moles thiosulfate.

Therefore, 1 mole thiosulfate = (1/6)*(moles of Cr2O72-)

The concentration of K2Cr2O7 = 0.049 mg/mL.

Molar mass of K2Cr2O7 = 294.185 g/mole.

Therefore, moles of K2Cr2O7 = (0.049 mg)*(1 g/1000 mg)*(1 mole/294.185 g) = 1.6656*10-7 mole.

Moles of thiosulfate neutralized = (1 mole K2Cr2O7)*(6 moles thiosulfate/1 mole K2Cr2O7) = 9.9936*10-7 mole.

Concentration of thiosulfate = (9.9936*10-7 mole)/[(1 mL)*(1 L/1000 mL)] = 9.9936*10-4 mol/L = 9.9936*10-4 M.

Volume of thiosulfate required for Se titration = 4.5 mL; therefore, moles of thiosulfate required for Se titration = (4.5 mL)*(9.9936*10-4 mol/L) = 4.49712*10-3 mmole.

As per stoichiometric equation (4),

Millimoles of I2 generated in Se titration = (4.49712*10-3 mmole)*(1 mole I2/2 moles thiosulfate) = 2.24856*10-3 mmole.

As per stoichiometric equation (3),

Millimoles of Se2O32- neutralized = millimoles of I2 generated = 2.24856*10-3 mmole.

As per stoichiometric equation (2),

2 mole SeBr4 = 1 mole Se2O32-

Therefore, millimoles of SeBr4 formed = (2.24856*10-3 mmole)*(2 mole SeBr4/1 mole Se2O32-) = 4.49715*10-3 mmole.

As per stoichiometric equation (1),

1 mole Se = 1 mole SeBr4.

Therefore, millimoles of Se in the sample = 4.49712*10-3 mmole.

Atomic mass of Se = 78.96 g/mole.

Therefore, mass of Se in the soil = (4.49712*10-3 mmole)*(1 mole/1000 mmole)*(78.96 g/1 mole)*(1000 mg/1 g) = 0.35509 mg.

Mass of soil taken = 10.0 g = (10.0 g)*(1 kg/1000 g) = 0.001 kg.

Therefore, concentration of Se in ppm = (0.35509 mg Se)/(0.001 kg soil) = 355.09 mg/kg ≈ 355.1 ppm (ans).